The difference between two numbers is ten. Four times the larger decreased by 30 equals three times the smaller increased by twenty. What are the numbers?
Answer by checkley77(12844)

x-y=10 or x=10+y
4x-30=3y+20
4(10+y)-30=3y+20
40+4y-30=3y+20
10+4y-3y=20
y=20-10
y=10 ans.
x-10=10
x=10+10
x=20 ans.
4*20-30=3*10+20
80-30=30+20
50=50